OpenStudy (anonymous):

How does 5k=2(mod 7) equal k=6(mod 7) ?

OpenStudy (anonymous):

k in an integer btw

OpenStudy (anonymous):

k is a*

OpenStudy (anonymous):

because if \[5k\equiv 2(\text{mod} 7)\] then also \[5k\equiv 2+4\times 7=30(\text{mod} 7)\] and so \[k\equiv 6(\text{mod} 7)\]

OpenStudy (anonymous):

well if u multiply the 2nd by 5, we get 5k = 30 (mod7)....but 30 (mod 7) is 2 (mod 7)

OpenStudy (anonymous):

also what qed said!

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

ty